# Yaroslav Aksenov, Ian Maksimov, Nikita Balagansky, Daniil Gavrilov Tinkoff

### v2

@@ Abstract
The complexity of the alignment problem stems from the fact that existing methods are considered unstable. Reinforcement Learning from Human Feedback (RLHF) addresses this issue by minimizing the KL divergence between the trained policy and the initial supervised fine-tuned policy (SFT) to avoid generating out-of-domain samples for the reward model (RM). Recently, many methods have emerged that shift from online to offline optimization, reformulating the RLHF objective and removing the reward model (DPO, IPO, KTO). Despite eliminating the reward model and the challenges it posed, these algorithms are still constrained in terms of closeness of the trained policy to the SFT one. In our paper, we argue that this implicit limitation in the offline optimization methods leads to suboptimal results. To address this issue, we propose a class of new methods called Trust Region (TR-DPO, TR-IPO, TR-KTO), which update the reference policy during training. With this straightforward update approach, we demonstrate the effectiveness of the new paradigm of language model alignment against the classical one on the Anthropic-HH and Reddit TL;DR datasets. Most notably, when automatically comparing TR methods and baselines side by side using pretrained Pythia 6.9B models on the Reddit TL;DR task, the difference in win rates reaches 8.4% for DPO, 14.3% for IPO, and 15% for KTO. Finally, by assessing model response ratings grounded on criteria such as coherence, correctness, helpfulness, and harmlessness, we demonstrate that our proposed methods significantly outperform existing techniques.
**footnotetext: Corresponding author: b.shaposhnikov@tinkoff.ai
@@ 1 Introduction
Figure 1: Comparison of the quality of the vanilla and Trust Region methods with soft and hard updates of the reference policy. For each of the methods (DPO, IPO, KTO), the win rate is presented in comparison with the SFT policy on the Anthropic HH dataset of the vanilla method (leftmost bar), a method in which the reference policy is updated according to the EMA principle with a coefficient of 0.6 (central bar), and a method with an update of the reference policy every 512 steps (rightmost bar). Pythia 2.8B models were used in the experiment. The comparison was carried out using GPT-4-0125-preview. See Section 2 for details of the Trust Region method and Section 4 for more experiments.
The alignment of Large Language Models (LLMs) is an increasingly pressing issue in contemporary Natural Language Processing (NLP). The primary goal is to train models that are not only effective, but also safe and controllable, which are qualities emphasized in recent research (Ouyang et al.,, 2022; Bai et al.,, 2022; Rafailov et al.,, 2023; Zhao et al.,, 2023). Achieving such safety typically involves fine-tuning LLMs to favor the generation of outputs that exhibit the desired behaviors.
@@
where $\mathcal{D}$ is the collection of training data, $\pi_{\theta}$ is the policy being optimized, $\pi_{\text{ref}}$ is the reference model (usually a supervised fine-tuned LM with the SFT policy), and $r_{\phi}(x,y)$ is the Reward Model (RM) trained in line with human preferences (Bradley and Terry,, 1952).
Initial attempts to address the issue of language model alignment employed Reinforcement Learning (RL) methods, where an RM, informed by human preferences, was developed. Subsequently, the LLM was tuned to produce outcomes aimed at maximizing the RM’s values (Bai et al.,, 2022), (Schulman et al.,, 2017). The current methodology has evolved to include a more intricate reparametrization of this procedure. For example, Direct Preference Optimization (DPO) by Rafailov et al., (2023) dispenses with the step of training the RM and directly optimizes the LLM by maximizing the training data likelihood as per the following loss function:
@@
Identity Preference Optimization (IPO) by Azar et al., (2023) slightly reformulates the original optimization task and replaces maximization of the reward with maximization of the probability that one text is better than the other. As a result, they obtain a different loss function:
@@
Ethayarajh et al., (2024) enhances the DPO method by adopting a Kahneman and Tversky, (1979) principle that losses outweigh equivalent gains. The Kahneman-Tversky Optimization (KTO) loss function can be defined as:
@@
where $z_{\text{ref}}=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\big{[}\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\big{]})$, $\lambda_{w}$ and $\lambda_{l}$ are coefficients controlling the degree of loss aversion (Kahneman and Tversky,, 1979).
This practice prompts us to question why the reference model remains static during training. For instance, consider a model aligned using a dataset $\mathcal{D}_{1}$ with a given reference policy. Then we collect more data, $\mathcal{D}_{2}$, which includes human preferences. The DPO approach suggests that, for further alignment with $\mathcal{D}_{2}$, the same reference model from the $\mathcal{D}_{1}$ training should be used (i.e., the SFT policy), even though the updated policy may now provide a more apt reference model. The same logic holds for IPO and KTO.
With this in mind, we introduce a novel concept to the training process for alignment algorithms, called the Trust Region (TR) approach. This includes methods such as TR-DPO, TR-IPO, and TR-KTO, where the reference policy is dynamically updated during training. This can be implemented either by softly integrating $\pi_{\theta}$ into $\pi_{\text{ref}}$ using a weighted approach, or by outright replacing the reference policy with $\pi_{\theta}$ after a predetermined number of steps. Our work’s contributions are as follows:
  • We introduce a novel approach for learning Preference Optimization using the Trust Region (TR) approach. This method outperforms the traditional ones, such as DPO, IPO, and KTO. Specifically, in model pair comparisons using GPT-4, TR-DPO with an $\alpha$ of 0.6 shows a 10% higher win rate than DPO for the Pythia 2.8B model. In the relative comparison of methods with SFT, the results of which are presented in Figure 1, TR-DPO with $\alpha$ of 0.6 shows a win rate 3% higher than DPO in comparison with the same SFT policy. This method’s efficacy is tested across two natural language generation tasks and three different model sizes.
  • Employing the Probability of Improvement (PoI) (Agarwal et al.,, 2021) metric, we demonstrate that TR methods statistically surpass the baselines in key human-centric (HC) metrics such as coherence, correctness, helpfulness, and harmlessness, validated on the Anthropic-HH dataset.
  • We analyze the dependency of HC metrics, KL divergence, and response diversity on the $\beta$ parameter in TR and classical methods. This analysis shows that policies trained by both our methods and classical ones diverge from the original SFT policy as $\beta$ decreases, yet the proposed methods demonstrate higher values of HC metrics at equivalent values of KL divergence.
@@ 2 Related Work
The alignment process is crucial for creating chat assistants and improving user satisfaction by training the model to generate safe, helpful, and correct responses (Bai et al.,, 2022). A fundamental method in adapting language models, Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al.,, 2020), has played an instrumental role in the success of groundbreaking LLMs such as GPT-3.5 (Ouyang et al.,, 2022; Ye et al.,, 2023), GPT-4 (OpenAI et al.,, 2024), and Llama 2 (Touvron et al.,, 2023). This method consists of three stages: Supervised Fine-Tuned Model (SFT Model) training, Reward Model (RM) training, and the Proximal Policy Optimization (PPO) algorithm (Schulman et al.,, 2017). PPO is used to train the model, or “policy” if following RL terminology, initialized by the SFT model, to maximize the expected reward of the model’s responses and reduce the Kullback-Leibler (KL) divergence between the trained policy and the SFT policy, as specified in Equation 1.
However, RLHF is not without its downsides, such as hyperparameter sensitivity, instability, implementation complexity, and high resource demand. Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) solved some problems by redefining the optimization problem and expressing a loss function to train the model directly, without RL algorithms. Similarly, Sequence Likelihood Calibration (SLiC) (Zhao et al.,, 2023) suggested using a margin loss between the probabilities of $y_{w}$ and $y_{l}$, which are responses generated by the SFT model. An extension of these methods was proposed with $\Psi$-PO (Azar et al.,, 2023), in which the focus shifted from maximizing reward to maximizing a certain probability function that $y_{w}$ is better than $y_{l}$. It was demonstrated that a specific case of this method, called IPO, could be more stable and resistant to overfitting than DPO. Another significant alignment alternative, Kahneman-Tversky Optimization (KTO) (Ethayarajh et al.,, 2024), seeks to accentuate human utility rather than solely relying on preference data. Rejection Sampling Optimization (RSO) incorporates features of both DPO and SLiC, suggesting that samples can be received from the optimal policy using rejection sampling. They can then be labeled by the reward model, and subsequently applied to train the policy using diverse loss functions (Liu et al., 2024a, ).
In the mentioned studies and the work of Wang et al., (2024), it was noticed that the term ensuring the trained policy’s closeness to the SFT policy does not significantly impact the final metrics. These observations spurred our idea to eliminate this limitation in the widely used DPO algorithm. Our proposed strategy involves updating the reference policy in a manner similar to updating the target network in value-based reinforcement learning algorithms (Awheda and Schwartz,, 2013). The following section delves deeper into how this has been accomplished.
3 Methods
@@
Figure 2: Schematic illustration of the proposed method. While vanilla DPO (left) uses a fixed reference policy during the training, for TR-DPO, we update it either with soft-update (center), for which parameters of $\pi_{\theta}$ are merged into parameters of $\pi_{\text{ref}}$ with some weight $\alpha$, or with hard-update (right), for which we copy parameters of $\pi_{\text{ref}}$ into a reference policy once, in a pre-determined number of training steps. See Section 2 for more details.
The alignment objective from Equation 1, and thus the objectives of offline optimization methods from Equations 2, 3 and 4, implies having a regularization with a fixed reference model for training. This objective involves maximizing the reward without moving far from the reference model. However, this requirement may seem too synthetic (e.g., see an example from Section 1), which is why there is a desire to move from a static reference model to an updated one during training.
@@
$\pi_{\text{ref}}\leftarrow\alpha\pi_{\theta}+(1-\alpha)\pi_{\text{ref}_{\text{prev}}},$ (5)
where $\alpha\in[0,1]$ is a weighting factor determining the rate at which the updates influence the reference policy. Since both $\pi_{\theta}$ and $\pi_{\text{ref}}$ are initialized from the same set of parameters, performing a soft update is justified by Rofin et al., (2022); Ilharco et al., (2023).
@@
$\pi_{\text{ref}}\leftarrow\pi_{\theta},$ (6)
which indicates a direct substitution of the reference model with the updated policy after a specified number of training iterations (denoted by $\tau\in\mathbb{N}$). This method provides more significant and episodic adjustments to the reference policy, promoting larger jumps in the model’s learning trajectory.
Policy updates can be applied to any LM alignment methods that maintain an implicit constraint on closeness to the reference policy. In this work, we experiment with the three most popular methods possessing the above-mentioned property: DPO (Rafailov et al.,, 2023), IPO (Azar et al.,, 2023), and KTO (Ethayarajh et al.,, 2024). We then propose a new class of methods called Trust Region (TR) methods: TR-DPO, TR-IPO, TR-KTO.
While the proposed change is straightforward to implement, updating the reference policy raises the question of how it changes the training objective. One way to think about it is to derive a connection to TR optimization methods (Schulman et al.,, 2015, 2017). Consider the following:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\Big{[}r_{\phi}(x,y)\Big{]}\text{, s.t. }\mathbb{D}_{\text{KL}}\Big{[}\pi_{\theta}(x,y)||\pi_{\theta_{\text{prev}}}(x,y)\Big{]}\leq\delta,$ (7)
where $\pi_{\theta_{\text{prev}}}$ is a previously obtained policy. Naive optimization with such an objective will lead to degeneration of the policy (Schulman et al.,, 2017). However, TR methods could be seen as approaches between the vanilla objective from Equation 1 and the TR variant from Equation 7, since we can control the frequency of changes in the reference policy. Such a connection explains the selection of TR parameters. A small $\alpha$ (large $\tau$) leads to rare reference policy updates, making the trained policy remain in the region close to the reference policy without large updates. This is equivalent to vanilla training with $\alpha=0$ or $\tau=\infty$. In contrast, a large $\alpha$ (small $\tau$) leads to frequent updates, making it possible for a policy to move far away from the reference one, as measured by $\mathbb{D}_{\text{KL}}$. It is similar to a pure TR setup with $\alpha=1$ or $\tau=1$.

### v3

# Learn Your Reference Model for Real Good Alignment

@@ Abstract
Despite the fact that offline methods for Large Language Models (LLMs) alignment do not require a direct reward model, they remain susceptible to overoptimization. This issue arises when the trained model deviates excessively from the reference policy, leading to a decrease in sample quality. We propose a new paradigm of offline alignment methods, called Trust Region (including variants TR-DPO, TR-IPO, TR-KTO), which dynamically updates the reference policy throughout the training process. Our results show that TR alignment methods effectively mitigate overoptimization, enabling models to maintain strong performance even when substantially deviating from the initial reference policy. We demonstrate the efficacy of these approaches not only through toy examples that exhibit reduced overoptimization, but also through direct, side-by-side comparisons in specific tasks such as helpful and harmless dialogue, as well as summarization, where they surpass conventional methods. Additionally, we report significant improvements in general-purpose assistant setups with the Llama3 model on the AlpacaEval 2 and Arena-Hard benchmarks, highlighting the advantages of Trust Region methods over classical approaches.
**footnotetext: Corresponding author: b.shaposhnikov@tbank.ru
@@ 1 Introduction
Aligning Large Language Models (LLMs) is an increasingly pressing issue in contemporary Natural Language Processing (NLP). The primary goal is to train models that are not only effective but also safe and controllable, qualities emphasized in recent research (Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2023; Zhao et al., 2023). Achieving such safety typically involves fine-tuning LLMs to favor the generation of outputs that exhibit desired behaviors.
@@
where $\mathcal{D}$ is the collection of training data, $\pi_{\theta}$ is the policy being optimized, $\pi_{\text{ref}}$ is the reference model (usually a supervised fine-tuned (SFT) policy), and $r_{\phi}(x,y)$ is the Reward Model (RM) trained to align with human preferences (Bradley & Terry, 1952).
Initial attempts to address the issue of language model alignment employed Reinforcement Learning (RL) methods, where an RM informed by human preferences was developed. Subsequently, the LLM was tuned to produce outputs that maximize the RM’s values (Bai et al., 2022; Schulman et al., 2017). Methods that use data generated from the policy during optimization are referred to as online alignment methods.
(a) AlpacaEval 2 (LC)
(b) ArenaHard
Figure 1: Evaluation performance of models trained by different methods, measured on the Alpaca Eval (a) and Arena Hard (b) benchmarks. The Llama-3-Base model was used as the baseline. The SFT stage was conducted on the UltraChat dataset, and the alignment stage on UltraFeedback. We compare vanilla methods (DPO, IPO, KTO) (left bars), their versions with a soft reference policy update (center bars), and with a hard update (right bars). Standard deviations are shown in the left image, while the 95% confidence intervals are indicated in the right one. See Section 4.3 for more details.
Current methodologies have evolved to include more intricate reparameterizations of this procedure, allowing us to omit sampling from the policy during training and instead optimize the model using pre-constructed datasets (i.e., offline alignment methods). For example, Direct Preference Optimization (DPO) by Rafailov et al. (2023) dispenses with the step of training the RM and directly optimizes the LLM by maximizing the likelihood of the training data, using the following loss function:
@@
Identity Preference Optimization (IPO) by Azar et al. (2023) slightly reformulates the original optimization task and replaces maximization of the reward with maximization of the probability that one text is better than the other. As a result, they obtain a different loss function:
@@
Ethayarajh et al. (2024) enhance the DPO method by adopting a Kahneman & Tversky (1979) principle that losses outweigh equivalent gains. The Kahneman-Tversky Optimization (KTO) loss function can be defined as:
@@
where $z_{\text{ref}}=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\big{[}\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\big{]})$, and $\lambda_{w}$ and $\lambda_{l}$ are coefficients controlling the degree of loss aversion (Kahneman & Tversky, 1979).
A key problem in aligning an LLM is reward overoptimization (Gao et al., 2022). Essentially, reward overoptimization occurs when the quality of the trained model decreases as the policy deviates from the reference policy (measured by KL divergence). This issue is commonly associated with reward hacking due to imperfections in the trained reward model. However, a similar pattern is observed in offline alignment methods, even though no explicit reward is used during training. Rafailov et al. (2024a) attribute overoptimization in offline methods to the inevitable increasing probability of Out-of-Domain (OOD) data.
In this paper, we show that overoptimization in offline alignment methods can be reduced by updating the reference policy during training, using what we call the Trust Region (TR) approach (as it resembles Trust Region optimization methods). This can be implemented either by softly integrating $\pi_{\theta}$ into $\pi_{\text{ref}}$ using a weighted approach or by directly replacing the reference policy with $\pi_{\theta}$ after a predetermined number of steps. Our work’s contributions are as follows:
  • We introduce Trust Region Alignment methods (TR-DPO, TR-IPO, TR-KTO) by incorporating reference policy updates through soft and hard update strategies, enhancing existing alignment techniques. We associate updates of the reference policy during training with overoptimization in offline alignment methods and confirm this with a toy MDP example.
  • Through extensive experiments, we demonstrate that our TR methods outperform their base counterparts across various model sizes on both task-specific and general benchmarks. For example, using pre-trained Pythia 6.9B models on the task-specific Reddit TL;DR summarization task, our methods achieve win rate improvements of 8.4% for DPO, 14.3% for IPO, and 15% for KTO over the baselines. Similarly, on the AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard (Li et al., 2024) general benchmarks with Llama3 (AI@Meta, 2024), our TR methods show significant win rate gains, with improvements of 10.8 points for DPO, 10.5 for IPO, and 5.6 for KTO compared to the classic methods (see Figure 1(b)).
  • We show that TR alignment methods reduce overoptimization by analyzing KL divergence and Human-Centric (HC) metrics, enabling the trained models to maintain better performance even as they diverge from the initial reference policy. At the same level of KL divergence, TR methods consistently achieve higher HC metrics compared to their classical counterparts.
@@ 2 Related Work
Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), based on PPO (Schulman et al., 2017), forms the foundation of language model alignment. Offline methods like DPO, IPO, KTO, RSO, and SimPO (Rafailov et al., 2023; Azar et al., 2023; Kahneman & Tversky, 1979; Liu et al., 2024b; Meng et al., 2024) simplify alignment further but suffer from reward overoptimization issues, as do online algorithms.
Overoptimization, well-studied in online methods (Gao et al., 2022; Coste et al., 2024), reveals that increasing KL divergence from the reference policy initially improves but eventually degrades policy quality. Similar patterns occur in offline methods without explicit reward models (Rafailov et al., 2024a), possibly due to increasing the probability of Out-of-Distribution (OOD) data. This problem remains unsolved in both approaches.
Notably, Wang et al. (2024) found that greater divergence from the reference policy does not always worsen results. This suggests that eliminating overoptimization might allow improvements by moving further from the reference model. Game-theoretical approaches (Rosset et al., 2024; Munos et al., 2024) have developed methods based on Nash equilibrium objectives, which imply updating reference policies during training. However, these inherently online methods rely on sampling from the policy and cannot be directly applied to the offline setup.
(a) DPO
[…]
3 Trust Region Alignment
3.1 Motivation
Following Rafailov et al. (2024b), during the training of offline alignment methods, the model moves away from the reference policy, thereby increasing the KL divergence between the resulting model and the reference model:
$\mathbb{D}_{\text{KL}}\Big{[}\pi_{\text{ref}}||\pi_{\theta}\Big{]}=-\mathbb{E}_{y\sim\pi_{\text{ref}(y|x)}}\Big{[}\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\Big{]},$ (5)
which inevitably decreases the average implicit reward of DPO. Since the reference policy is usually trained with chosen sequences, Rafailov et al. (2024b) hypothesize that these chosen sequences must reduce their probability during optimization. Because $\pi_{\theta}$ is a probability distribution, reducing the probabilities of some sequences under it (in practice, reducing the probabilities of in-domain (ID) data) leads to increasing the probabilities of other sequences (e.g., out-of-domain (OOD) data), which is associated with overoptimization. Notably, when considering the gradient of the DPO objective:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=\\ &=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right],\end{split}$ (6)
it is unclear why such behavior occurs. The DPO objective should force $\log\frac{{\pi_{\theta}(y_{w}|x)}}{{\pi_{\theta}(y_{l}|x)}}$ to be larger than $\log\frac{{\pi_{\text{ref}}(y_{w}|x)}}{{\pi_{\text{ref}}(y_{l}|x)}}$. This could potentially be achieved by gradually increasing the probabilities of chosen sequences and significantly decreasing the probabilities of rejected sequences, thereby decreasing the average implicit reward without shifting probability mass to OOD examples. However, in practice, both of these probabilities usually decrease over the course of training (Pal et al., 2024).
From this perspective, the dynamics of the change in probabilities of sequences are crucial when discussing overoptimization. Consider the Hessian of the DPO objective:
$\nabla^{2}_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma(s)\nabla^{2}_{\theta}s+\sigma(s)\big{(}1-\sigma(s)\big{)}\nabla_{\theta}s\big{(}\nabla_{\theta}s\big{)}^{\top}\right],$ (7)
where $s=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ (see Appendix Section A for derivation). During training, $s$ usually moves from 00 to smaller values, causing $\sigma(s)$ to approach 00. This leads to a vanishing curvature in the loss landscape. We hypothesize that due to this effect, if we initially enter a phase where the probabilities of ID data are decreasing, it becomes difficult to reverse this process, potentially leading to overoptimization.
To mitigate this behavior, we propose updating the reference policy $\pi_{\text{ref}}$ during training, for example, by occasionally setting $\pi_{\text{ref}}\leftarrow\pi_{\theta}$. This approach allows us to "reset" the optimization process and increase the curvature of the loss landscape, possibly preventing us from getting stuck in a cycle of reducing the probabilities of chosen sequences. By doing so, we relax the initial constraints on divergence from the SFT policy. If our hypothesis on reducing overoptimization by doing this is correct, then such relaxation is not harmful. Moving away from the SFT policy is not inherently bad; if it leads to improved metrics, it is beneficial.
@@ 3.1 Motivation
Figure 3: Schematic illustration of the proposed method. While vanilla DPO (left) uses a fixed reference policy during the training, for TR-DPO, we update it either with soft-update (center), for which parameters of $\pi_{\theta}$ are merged into parameters of $\pi_{\text{ref}}$ with some weight $\alpha$, or with hard-update (right), for which we copy parameters of $\pi_{\theta}$ into a reference policy once, in a predetermined number of training steps. See Section 3.2 for more details.
Notably, this process resembles Trust Region (TR) optimization methods, where we optimize a function within a predefined range. In our case, for each step $i$, this range is defined by the reference policy $\pi_{\text{ref}}^{(i-1)}$. This logic can also be applied to other offline alignment methods like IPO or KTO. Therefore, it is interesting to consider whether updating the reference policy using these methods could be beneficial.
In the following section, we describe a practical implementation of updating the reference model.
3.2 Method
@@
$\pi_{\text{ref}}\xleftarrow{S.G.}\alpha\pi_{\theta}+(1-\alpha)\pi_{\text{ref}_{\text{prev}}},$ (8)
where $S.G.$ denotes the stop-gradient operation, and $\alpha\in[0,1]$ is a weighting factor that determines the rate at which the updates influence the reference policy. This update is performed at each optimization step, thereby softly updating the reference policy. Since both $\pi_{\theta}$ and $\pi_{\text{ref}}$ are initialized from the same set of parameters, performing a soft update is justified by Rofin et al. (2022); Ilharco et al. (2023).
@@
$\pi_{\text{ref}}\xleftarrow{S.G.}\pi_{\theta},$ (9)
which indicates a direct substitution of the reference model with the updated policy after a specified number of training iterations $\tau\in\mathbb{N}$. This method provides more significant and episodic adjustments to the reference policy, promoting larger jumps in the model’s learning trajectory.
Both methods involve fully replacing the reference policy’s parameters, and they are not directly optimized through gradient updates. Reference policy updates can be applied to any LM alignment methods that maintain an implicit constraint on closeness to the reference policy. In this work, we experiment with the three most popular methods possessing the above-mentioned property: DPO (Rafailov et al., 2023), IPO (Azar et al., 2023), and KTO (Ethayarajh et al., 2024). We then propose a new class of methods called Trust Region (TR) methods: TR-DPO, TR-IPO, TR-KTO.
Following Rafailov et al. (2024a), we experimented with a toy MDP example (see Appendix Section B for more details), for which we trained the proposed methods and compared them with their vanilla variants. See Figure 2 for a comparison of DPO with TR-DPO, and Appendix Figures 8, 9 for IPO/TR-IPO and KTO/TR-KTO, respectively. DPO and IPO show an increase in the probabilities of OOD examples, while TR-DPO and TR-IPO increase the probabilities of chosen trajectories and decrease the probabilities of OOD trajectories. These results support our initial hypothesis on updating the reference policy. In this toy example, we do not see signs of overoptimization with the toy MDP setup for the vanilla KTO method; therefore, in the following sections, we experimentally demonstrate that all of our proposed methods show higher quality on real tasks, which can be attributed to their reduced tendency for overoptimization.